Solve. 4(x + 2) – 3x = x – 6
Answer by fractalier(2101)

Okay, step-by-step we get
4(x + 2) – 3x = x – 6
4x + 8 - 3x = x - 6
x + 8 = x - 6
8 = -6
When does occur?
Never.
Therefore there is no solution to this problem.